OpenStudy (anonymous):

Using complete sentences, explain what the discriminant is and what it tells you about the solutions of a quadratic equation. Provide a unique example to back up your explanation.

OpenStudy (karatechopper):

hello

OpenStudy (across):

Every 2nd order polynomial has three possible, general solutions: two unique roots, one repeated root and complex conjugate roots. The discriminant of a polynomial helps you determine which of those three cases you have.

OpenStudy (paxpolaris):

\[When\ ax^2+bx+c=0\]\[x={-b \pm \sqrt{b^2-4ac}\over2a}\]

OpenStudy (across):

For example, take\[x^2+2x-3.\]Its discriminant will be\[\sqrt{2^2-4(1)(-3)}=\sqrt{4+12}=\sqrt{16}=4.\]Since \(4>0\), you know that this polynomial has two unique solutions.

OpenStudy (paxpolaris):

The discriminant is:\[b^2-4ac\]

OpenStudy (across):

For example, take\[x^2+4x+4.\]Its discriminant will be\[\sqrt{4^2-4(1)(4)}=\sqrt{16-16}=\sqrt{0}=0.\]Since your discriminant is \(0\), you know that your polynomial will have a repeated root.

OpenStudy (anonymous):

thanks guys great answers

OpenStudy (across):

For example, take\[x^2+x+1.\]Its discriminant will be\[\sqrt{1^2-4(1)(1)}=\sqrt{1-4}=\sqrt{-3}.\]Since its (correct)* discriminant will be \(-3\) and since that is less than zero, you know that your polynomial will have complex roots. * I say correct because I took the square root of the discriminant in the above examples.

OpenStudy (across):

Forget the square root when computing discriminants.